OpenStudy (anonymous):

What is the limit of (n!+2^n)^(1/n) as n->infinity?

OpenStudy (btaylor):

just to check...\[\lim_{n \rightarrow \infty} (n!+2^n)^\frac{1}{n}\]Is that right?

OpenStudy (anonymous):

Indeed. I do not understand why the answer is infinity. I have tried to use the Squeeze Theorem, and got a finite value.. :/

OpenStudy (btaylor):

Huh...I can't think of anything. The inside argument (n!+2^n) would go to infinity very quickly

OpenStudy (anonymous):

Yes, but is there any mathematical solution for this, and not just intuition? :)

OpenStudy (btaylor):

Graphing it: http://awesomescreenshot.com/00a2fffj55

OpenStudy (anonymous):

Even on a test? :D (this question was taken from a test)

OpenStudy (btaylor):

Hmmm...(2^x)^(1/x) turns into just 2 (except at x=0) I don't really know

OpenStudy (anonymous):

Because the factorial is the most powerful operation. So, you only need to pay attention to what it does as it goes to infinity

OpenStudy (btaylor):

@donny471 what does it mean that factorial is most powerful?

OpenStudy (anonymous):

Just as exponentials have more of an effect than polynomials, factorials dominate over both of them

OpenStudy (anonymous):

So, we basically ignore the other operations since they offer a minimal influence on the overall behavior of the function

OpenStudy (anonymous):

Exponentials are the fastest growing to infinity.. way faster than factorials..

OpenStudy (anonymous):

Think of 2^10 vs 10!, even for a low n which is 10, the logarithem is much much bigger..

OpenStudy (anonymous):

10! is drastically larger than 2^10

OpenStudy (anonymous):

OMG, MY BAD -_-

OpenStudy (anonymous):

If we arrange things, which is the fastest growing to infinity? (and the slowest).. I know that ln is really slow, but do you know any other important ones? 1. x^n (x>0) 2. n! 3. ln(n) 4. n^n

OpenStudy (anonymous):

(i just listed a few, the order has no meaning)

OpenStudy (anonymous):

the n^n is most powerful there

OpenStudy (anonymous):

If you're ever uncertain just plug values into a calculator

OpenStudy (anonymous):

lim of  as n->infinity. Why does the factorial lose here?

OpenStudy (anonymous):

plug in 2 for n to start. Then 20... Whichever one overflows your calculator first is more powerful.

OpenStudy (anonymous):

Really it's whichever one grows faster. If you have a dingy calculator it'll overflow quickly though

OpenStudy (anonymous):

the problem is that we are not allowed to use calculators in test.. and this one is taken from a test :/

OpenStudy (anonymous):

Use Sterling's formula to deduce that \[ n!\approx \left(\frac{n}{e}\right)^n \sqrt{2 n \pi }>\left(\frac{n}{e}\right)^ n \\Hence\\ (n!)^{1/n}>\frac{n}{e}\\ \lim_{n\to \infty } (n!)^{1/n} > \lim_{n\to \infty }\frac{n}{e}=\infty\\ \]

OpenStudy (anonymous):

Your limit is bigger than the above. So you are done. See Stirling's formula on http://en.wikipedia.org/wiki/Stirling%27s_approximation

OpenStudy (anonymous):

Thanks eliassaab!

OpenStudy (anonymous):

YW